A 0.100 M solution of a weak acid, HX, has a pH of 2.610. What is the value of Ka for HX?

1)6.03 x 10 -6

2)6.03 x 10 -4

3)2.45 x 10 -2

4)2.45 x 10 -1

5)6.03 x 10 -5

6)2.45 x 10 -3



1
Expert's answer
2021-06-07T03:15:32-0400

The K

a

 expression is:

Ka=[A-M-][H3O)O+][HA]

=x×x0.100−x

∴ Ka=x2

0.100−x

=0.100^2

= 2.45 × 102